Take any number (except for 1). Square that number and then subtract one. Divide by one less than your original number. Now subtract your original number. Did you reached 1 for an answer? You should have. How does this number game work? (Hint: Redo the number game using a variable instead of an actual number and rewrite the problem as one rational expression). How did the number game use the skill of simplifying rational expressions? Create your own number game using the rules of algebra and post it for your classmates to solve. Be sure to think about values that may not work. State whether your number game uses the skill of simplifying rational expressions. : Take any number (except for 1). Square that number and then subtract one. Divide by one less than your original number. Now subtract your original number. Did you reached 1 for an answer? You should have. How does this number game work? (Hint: Redo the number game using a variable instead of an actual number and rewrite the problem as one rational expression). How did the number game use the skill of simplifying rational expressions? Create your own number game using the rules of algebra and post it for your classmates to solve. Be sure to think about values that may not work. State whether your number game uses the skill of simplifying rational expressions.
Answer by oscargut(666)

x=original number
Square that number and then subtract one. so we have x^2-1
Divide by one less than your original number then we have (x^2-1)/(x-1)
Now subtract your original number so the result is (x^2-1)/(x-1)-x
=(x+1)(x-1)/(x-1) - 1 = (x+1)-x = 1
so the result is always 1 but the original number is not 1 because we can not divide by x-1(x-1 is 0 when x=1)